OpenStudy (anonymous):

solve the equation 2x-4y=16 for y and find the value of y when x=-1,0, and 3 solve the distance traveled formula for r

OpenStudy (anonymous):

2x-4y=16 -2x -2x -4y=-2x+16 divide each side by -4 y=1/2x-4

OpenStudy (anonymous):

thnx!

OpenStudy (anonymous):

To find the value of y when equals the numbers you put down, all you have to do is plug it in to the equation... For example: y=1/4(-1)-4

OpenStudy (anonymous):

no prob!

OpenStudy (anonymous):

i have more if u wanna help me

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I'll explain it to you this time so you can understand it better and so you're able to do it on your own

OpenStudy (anonymous):

Just give me the equations

OpenStudy (anonymous):

give an example of an open,true,and false statement explain how to determinr if a is a solution to an equation make up your own word problem and solve it

OpenStudy (anonymous):

hang on...

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

open: 5x+8=3(x+5) true: 5+3=8 false: 3+3=7

OpenStudy (anonymous):

thank you can u help me with the word problem?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

could you wait a couple minutes again?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Ok i'm back

OpenStudy (anonymous):

wait, so do you need me to determine if a is a solution to the problem?

OpenStudy (anonymous):

or just make a word problem?

OpenStudy (anonymous):

make a word problem and solve it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Wait have you learned how to solve systems yet?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so here we go: Sally has to buy some gift baskets. Each costs $15. She has $180 to spend. Write an equation and solve for x.

OpenStudy (anonymous):

Here's the equation: 180=15x

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (anonymous):

No prob! Do you get how I did everything?

OpenStudy (anonymous):

yes im understanding. please friend me so i can inbox you if i ever need help!?

OpenStudy (anonymous):

Sure! Could you give me a medal for my answer?

OpenStudy (anonymous):

yupp ofcorse!

OpenStudy (anonymous):

Thanks!